OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9.

OpenStudy (anonymous):

wait, x=-11

OpenStudy (anonymous):

so its -11?

OpenStudy (anonymous):

yes -11

OpenStudy (anonymous):

Simplifying 5x + 2 = 4x + -9 Reorder the terms: 2 + 5x = 4x + -9 Reorder the terms: 2 + 5x = -9 + 4x Solving 2 + 5x = -9 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 2 + 5x + -4x = -9 + 4x + -4x Combine like terms: 5x + -4x = 1x 2 + 1x = -9 + 4x + -4x Combine like terms: 4x + -4x = 0 2 + 1x = -9 + 0 2 + 1x = -9 Add '-2' to each side of the equation. 2 + -2 + 1x = -9 + -2 Combine like terms: 2 + -2 = 0 0 + 1x = -9 + -2 1x = -9 + -2 Combine like terms: -9 + -2 = -11 1x = -11 Divide each side by '1'. x = -11 Simplifying x = -11